### v1

# Information Re-Organization Improves Reasoning in Large Language Models

Abstract
Improving the reasoning capabilities of large language models (LLMs) has attracted considerable interest. Recent approaches primarily focus on improving the reasoning process to yield a more precise final answer. However, in scenarios involving contextually aware reasoning, these methods neglect the importance of first identifying logical relationships from the context before proceeding with the reasoning. This oversight could lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning outcomes. In this paper, we propose an information re-organization (InfoRE) method before proceeding with the reasoning to enhance the reasoning ability of LLMs. We first perform a re-organization processing of the contextual content, e.g., documents or paragraphs, to obtain logical relationships. Then, we utilize the re-organized information in the reasoning process. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships. To demonstrate the effectiveness of our approach in improving the reasoning ability, we conduct experiments using Llama2-70B, GPT-3.5, and GPT-4 on various contextually aware multi-hop reasoning tasks. Using only a zero-shot setting, our method achieves an average improvement of 3% across all tasks, highlighting its potential to improve the reasoning performance of LLMs. Our source code is available at https://github.com/hustcxx/InfoRE.
22footnotetext:   Corresponding author.
1 Introduction
Large language models (LLMs) demonstrate powerful generative capabilities and achieve remarkable performance across a range of linguistic tasks Brown et al. (2020); Touvron and et al. (2023); OpenAI (2023). However, their capabilities in performing complex reasoning tasks — an essential aspect of advanced language understanding and intelligent decision-making — still present substantial challenges Arkoudas (2023); Blair-Stanek et al. (2023). This has spurred researchers to explore innovative strategies Wei et al. (2022); Yao et al. (2023); Besta et al. (2023) to improve the reasoning capabilities of these models.
Figure 1: InfoRE (Ours) vs existing methods for improving the reasoning of LLMs. In contrast to the existing methods that primarily focus on the reasoning process, our InfoRE emphasizes the re-organization of context information. The [TEXT] in italics indicates that it is optional in the reasoning process.
Recently, diverse methods have been developed to enhance the reasoning ability of LLMs. For example, a notable method Chain-of-Thought (CoT) Wei et al. (2022), incorporates a series of intermediate reasoning steps into the reasoning. CoT Wei et al. (2022) allows for a more transparent and understandable path to the final answer, making it easier to follow the logic behind the conclusion. Building upon this foundation, subsequent approaches such as Tree of Thoughts (ToT) Besta et al. (2023) and Graph of Thoughts (GoT) Besta et al. (2023) are proposed to further refine the reasoning steps and enhance the accuracy and reliability of LLMs. Different from the sequential intermediate steps of Chain-of-Thought Wei et al. (2022), Tree of Thoughts Besta et al. (2023) and Graph of Thoughts Besta et al. (2023) model the problem solving process into structured tools of tree and graph, respectively. A critical observation is that these existing approaches primarily focus on improving the reasoning process of large language models, as shown in Figure 1 (left). However, in scenarios involving contextually aware reasoning, it is equally important to first identify logical relationships of context before proceeding with the reasoning, not just improve their reasoning process. This is because logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of reasoning Braine (1978). Nevertheless, these existing methods often neglect this crucial step. Such an oversight can lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning results.
Inspired by the fact that when faced with context-aware reasoning tasks humans often first re-organize existing contextual information to uncover the logical relationships and enhance their understanding of the context, we propose an information re-organization (InfoRE) method, to ground reasoning by the re-organized information. As shown in Figure 1 (right), different from previous methodologies that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, our approach takes a novel direction of context re-organization. We emphasize the utilization of re-organized contextual content to explicitly present the logical relationships that are often implicit within the plain text, promoting more effective reasoning. Specifically, we first perform a re-organization processing of the contextual content into a MindMap structure Buzan et al. (2010), uncovering logical relationships implicit within the context. We employ this structure for re-organized information because it is rich in logical relationships and encompasses multi-hop connections. Then, we utilize the re-organized context to reason. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships, facilitating the quality and reliability of reasoning. Besides, our information re-organization method can be integrated with existing prompt methods, like CoT Wei et al. (2022), to further improve the reasoning ability of LLMs. To verify the efficacy of our proposed InfoRE method, we conduct experiments using Llama2-70B Touvron and et al. (2023), GPT-3.5 Brown et al. (2020), and GPT-4 OpenAI (2023) on various contextually aware multi-hop reasoning tasks, including claim verification Guo et al. (2021), question answering Perez et al. (2020), and reading comprehension Dua et al. (2019). Using only a zero-shot setting, our method achieves an average improvement of 3% across all tasks, highlighting its potential to improve the reasoning performance of LLMs.
Our main contributions are as follows:
  • In contrast to existing methodologies that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, we take a novel direction of context re-organization.
  • We first perform a re-organization processing of the contextual content to uncover the logical relationships that encompass multi-hop connections of the existing context, then to facilitate the reasoning of LLMs.
  • Experiment improvements on contextually aware multi-hop reasoning tasks across claim verification, question answering, and reading comprehension show the efficacy of our proposed method.
2 Related Work
Reasoning with LLMs
Large language models (LLMs) Touvron and et al. (2023); Brown et al. (2020); OpenAI (2023), have revolutionized the field of natural language processing (NLP) and demonstrate remarkable proficiency across a range of linguistic tasks. To further improve the reasoning ability of LLMs has attracted considerable interest. In-context learning Brown et al. (2020), as a promising approach to enhance the reasoning abilities of LLMs, has been verified in mathematical reasoning task Wei et al. (2022). In addition, Chain of Thought (CoT) Wei et al. (2022) incorporates a coherent series of intermediate reasoning steps to improve the reasoning ability of LLMs. Following this paradigm, Tree-of-Thoughts (ToT) Yao et al. (2023) and Graph-of-Thoughts (GoT) Besta et al. (2023) have also been proposed to focus on improving the structure of the intermediate chain. Then self-consistency Wang et al. (2022) and plan-to-solve Wang et al. (2022) focus on the reliability of the chain. Recently, the step-back prompting Zheng et al. (2023) is proposed, which obtains the high-level concept and first principles from instances by abstraction in the first step, then guides the reasoning of LLMs with the obtained concept and principles. For tasks that require multi-hop reasoning, current approaches Pan et al. (2023); Radhakrishnan et al. (2023) generally decompose multi-hop problems into simpler sub-tasks problems relying on the in-context learning ability of LLMs. In contrast to existing methods, we take a novel direction of context re-organization to enhance the reasoning capabilities of LLMs for contextually aware reasoning tasks.
Information Re-organization
Information re-organization is a technique that leverages other structures to improve the clarity and comprehensibility of information. Moreover, information re-organization can also reveal conceptual relationships implicit in the original textual text. The graph, as the most common re-organization structure, is often used to enhance contextual representation in a variety of tasks. For example, BASS Wu et al. (2021) re-organizes the original document to a semantic graph to improve the summary generation. DGM Ouyang et al. (2021) re-organizes the context to an explicit discourse graph in the reading comprehension task. MindMap Buzan et al. (2010) is a powerful structure method for representing knowledge and concepts, it can be used to construct a hierarchical abstraction of natural language text Elhoseiny and Elgammal (2014). Previous method Dalamagas et al. (2010) uses it to organize a large amount of scientific material to enhance the clarity of the text material. In this paper, we use it as the structure of re-organized information to uncover the logical relationships and multi-hop connections implicit within the plain context. Then we use this re-organized information to further improve the reasoning ability of LLMs on multi-hop reasoning tasks.
3 Methodology
In this section, we first give a formulation of a context-aware reasoning task in §3.1 and then describe our method in detail. As shown in Figure 2, our method consists of two steps, an information re-organization step §3.2 and a reasoning step with the re-organized context information §3.3.
[…]
3.1 Task Formulation
Given a sample $(\mathcal{C},\mathcal{Q},\mathcal{A})$, where $\mathcal{C}$ denotes gold context, $\mathcal{Q}$ denotes question, the task aims to obtain answer $\mathcal{A}$ to the question $\mathcal{Q}$, based on gold context $\mathcal{C}$ with a large language model. In particular, it requires multi-hop reasoning to get the answer $\mathcal{A}$ to question $\mathcal{Q}$.
3.2 Information Re-Organization
For a question $\mathcal{Q}$, if the derivation of its answer $\mathcal{A}$ relies on context $\mathcal{C}$, then a deeper understanding of $\mathcal{C}$ is crucial. Logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of understanding and reasoning Braine (1978). However, the logical relationships in the original plain context are often implicit. Therefore, we uncover to obtain logical relationships in the context through information re-organization. MindMap Buzan et al. (2010) is a powerful structure for representing knowledge, concepts, and ideas, which can provide a hierarchical abstraction of plain text. Converting a text paragraph to this structure would provide an easier way to access the knowledge and ideas in the context. In addition, we choose this as the structure for re-organized information because it contains not only logical relationships but also multi-hop connections.
For example, for a multi-hop question: Where was the director of film The Green Fog born?, the corresponding context: The Green Fog is an experimental film directed by Guy Maddin that loosely revisits the plot of Alfred Hitchcock’s movie Vertigo through a collage of found footage … Guy Maddin (born February 28, 1956) is a Canadian screenwriter, director, author, cinematographer, and film editor of both features and short films, as well as an installation artist, from Winnipeg, Manitoba …, and the re-organized context is:
The Green Fog:
  Type: experimental film
  Director:
     Name: Guy Maddin
     Birth date: February 28, 1956
     Nationality: Canadian
     City: Winnipeg, Manitoba
     Profession: screenwriter, director
  plot: Alfred Hitchcock’s movie
[…]
Formally, MindMap elements are organized in hierarchical branches and reflect the logical relationships of elements. The same level can reflect parallel relationships and indented levels reflect causal relationships, as shown in Figure 2. In terms of content, MindMap is often around a central topic or idea. Based on the format and content criteria of the MindMap, we utilize the generative capabilities of large language models to construct it. According to the evaluation Chen et al. (2021), large language models are capable of generating information with a logical structure. Specifically, the large language model is asked to re-organize the context to a MindMap following instructions according to the context information and the question. Detailed information is shown in Appendix A.
3.3 Reasoning With Re-Organized Information
Following the re-organization process in Section 3.2, we get the re-organized context $\mathcal{G}$ from the context $\mathcal{C}$. Then the re-organized information $\mathcal{G}$ can be used as a context alone or combined with the original $\mathcal{C}$ to get the final answer of $\mathcal{Q}$. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships, facilitating the quality and reliability of reasoning.
Specifically, a large language model processes this context to infer the answer to question $\mathcal{Q}$ with a prompt. This step relies on the advanced natural language processing capabilities of large language models, encompassing understanding the context, integrating information, and logical reasoning based on this understanding. The output of a large language model is a response to question $\mathcal{Q}$, based on its comprehension of the context and analysis. The overall prompt template consists of three components: original context, e.g., documents or paragraphs, re-organized information, and a question. The prompt template is as follow:
Context:
   [Re-Organized TEXT]
   [TEXT][Optional]
Question:
   [QUESTION]
please answer the question based on the
Context.
This prompt can be combined with existing methods such as Chain-of-Thought (CoT) Wei et al. (2022) by appending “Let’s think step by step.” to the question to further improve the accuracy of reasoning. Then we use a regex expression to extract answer $\mathcal{A}$ from the response.
4 Experimental
4.1 Tasks and Datasets
To verify the effectiveness of our information re-organization method, we conduct experiments across a range of contextually aware multi-hop reasoning tasks and datasets, including claim verification Guo et al. (2021), question answering Perez et al. (2020), and reading comprehension Dua et al. (2019). The detailed dataset information we used is displayed in Table 1.
Task Dataset Split Pairs
Claim Verification FEVEROUS Dev 2959
HOVER Dev 4000
SCIFACT Test 212
Question Answering 2WMHopQA Dev 500
StrategyQA Dev 229
MusiQue Dev 2417
Reading Comprehension HotpotQA Dev 500
WIKIHOP Dev 500
Table 1: Details statics information of evaluation datasets we used in the paper. Pairs denote the number of examples. 2WMHopQA is the short name of 2WikiMultiHopQA.
Claim Verification
The task involves assessing a given claim against a set of evidence documents to determine whether they support or refute the claim Guo et al. (2021). We consider HOVER Jiang et al. (2020) and FEVEROUS Aly et al. (2021), which comprise complex claims that necessitate multi-hop reasoning for verification. Additionally, we also take into account the SCIFACT Wadden et al. (2022) dataset, notable for its inclusion of scientific claims.
LLMs Methods HOVER (2-hop) HOVER (3-hop) HOVER (4-hop) FEVEROUS SCIFACT
LLAMA2-70B Standard 49.41 48.35 47.82 63.39 60.70
InfoRE 51.66 50.30 49.45 67.01 63.39
$\uparrow 2.25$ $\uparrow 1.95$ $\uparrow 1.63$ $\uparrow 3.62$ $\uparrow 2.69$
CoT 50.02 48.76 48.01 64.53 61.24
InfoRE + CoT 52.40 51.21 50.07 67.96 64.24
$\uparrow 2.38$ $\uparrow 2.45$ $\uparrow 2.06$ $\uparrow 3.43$ $\uparrow 3.00$
GPT-3.5 Standard 66.77 66.10 63.91 92.00 77.42
InfoRE 69.10 67.03 64.39 94.00 80.65
$\uparrow 2.33$ $\uparrow 0.93$ $\uparrow 0.48$ $\uparrow 2.00$ $\uparrow 3.23$
CoT 68.66 66.78 64.16 92.53 75.27
InfoRE + CoT 70.24 67.86 65.16 94.32 81.72
$\uparrow 1.58$ $\uparrow 1.08$ $\uparrow 1.00$ $\uparrow 1.79$ $\uparrow 6.45$
GPT-4 Standard 72.40 71.02 70.06 92.33 91.40
InfoRE 75.42 73.95 71.22 94.67 93.55
$\uparrow 3.02$ $\uparrow 2.93$ $\uparrow 1.16$ $\uparrow 2.34$ $\uparrow 2.15$
CoT 73.82 72.07 70.68 92.67 92.47
InfoRE + CoT 75.87 74.26 72.08 95.64 94.77
$\uparrow 2.05$ $\uparrow 2.19$ $\uparrow 1.40$ $\uparrow 2.97$ $\uparrow 2.30$
Table 2: Zero-shot performance on claim verification task across three large language models.
Question Answering
For this task, we consider the following datasets: 2WikiMultiHopQA Ho et al. (2020), StrategyQA Geva et al. (2021), MuSiQue Trivedi et al. (2021), and HotpotQA Yang et al. (2018). To answer the questions in these datasets requires not only multi-hop reasoning but also cross-document analysis.
Reading Comprehension
Machine reading comprehension task requires a model to process documents and select an answer from the provided candidates to a question about the content Dua et al. (2019). We primarily consider WIKIHOP Welbl et al. (2017) in the task, which necessitates multi-hop reasoning to derive the final answer. Additionally, HotpotQA dataset Yang et al. (2018) is also frequently considered as part of the question answering domain.
4.2 Language Models and Baselines
Models
In our paper, we employ pre-trained large language models such as Llama2-70B Touvron and et al. (2023), GPT-3.5 (text-davinci-003) Ouyang et al. (2022) and GPT-4 OpenAI (2023) as the foundational models. We configure all models with top_p parameter as $1.0$ and temperature as $0.0$.
Baselines
1) Standard. 2) CoT. The Standard approach is a method that directly reasons with the original textual context. The Chain of Thought (CoT) Wei et al. (2022) method involves augmenting standard reasoning methods by adding a step-by-step thought process. In our paper, we conduct the CoT strategy by appending the sentence “Let’s think step by step.” at the end of the question.
In this paper, we conduct experiments, including both the baseline methods and our InfoRE in a zero-shot setting to counteract the potential randomness associated with demonstrations in a few-shot setting. We also design an answer-format instruction within the prompts for various tasks to standardize the structure of the final answer, thereby enhancing the precision of answer extraction. Moreover, all results reported in the paper use only the reorganized contextual information to reason. Comprehensive details about prompts and answer-format instruction are available in Appendix B.
4.3 Evaluation Metrics
We use F1 to measure the results on all the datasets we used in our paper following previous methods Pan et al. (2023); Radhakrishnan et al. (2023); Aly et al. (2021). We run the official evaluation scripts of each dataset to get the output scores.
LLMs Methods 2WikiMultiHopQA MuSiQue StrategyQA HotpotQA WIKIHOP
LLAMA2-70B Standard 52.56 49.55 51.23 66.07 40.32
InfoRE 56.83 52.25 54.86 69.73 42.02
$\uparrow 4.27$ $\uparrow 2.70$ $\uparrow 3.63$ $\uparrow 3.66$ $\uparrow 1.70$
CoT 52.99 52.90 56.80 66.80 41.07
InfoRE + CoT 57.67 56.07 59.36 70.32 42.77
$\uparrow 4.68$ $\uparrow 3.17$ $\uparrow 2.56$ $\uparrow 3.52$ $\uparrow 1.70$
GPT-3.5 Standard 58.25 55.01 59.39 73.30 48.92
InfoRE 63.80 57.38 62.78 76.53 51.13
$\uparrow 5.55$ $\uparrow 2.37$ $\uparrow 3.39$ $\uparrow 3.23$ $\uparrow 2.21$
CoT 59.37 57.05 67.51 73.90 49.65
InfoRE + CoT 64.51 59.87 69.87 77.20 52.05
$\uparrow 5.14$ $\uparrow 2.82$ $\uparrow 2.36$ $\uparrow 3.30$ $\uparrow 2.40$
GPT-4 Standard 72.69 62.65 68.32 79.33 55.46
InfoRE 76.27 66.02 70.46 82.81 57.66
$\uparrow 3.58$ $\uparrow 3.37$ $\uparrow 2.14$ $\uparrow 3.48$ $\uparrow 2.20$
CoT 74.08 64.36 68.50 80.66 56.02
InfoRE + CoT 78.43 68.82 70.88 83.75 58.32
$\uparrow 4.35$ $\uparrow 4.46$ $\uparrow 2.38$ $\uparrow 3.09$ $\uparrow 2.30$
Table 3: Zero-shot results on Question Answering and Reading Comprehension tasks.
5 Results
5.1 Claim Verification
Table 2 presents a comprehensive performance comparison between our InfoRE and existing zero-shot techniques. For the HOVER dataset, we segment it into the 2-hop, 3-hop, and 4-hop levels according to the hop number in the dataset following ProgramFC Pan et al. (2023). As depicted in Table 2, our InfoRE demonstrates significant improvements in the zero-shot claim verification task. The CoT Wei et al. (2022) approach offers a lightly increase of 0.62% in the HOVER 4-hop using GPT-4, which indicates its marginal utility in more complex reasoning scenarios. Yet, our InfoRE achieves 1.16% improvement on the HOVER 4-hop using GPT-4 showing remarkable performance on contextual aware understanding and reasoning. This improvement is further increased to 72.08% in combination with CoT, suggesting that the methods complement each other effectively.
In the case of Llama2-70B, the combined application of InfoRE and CoT yields a score of 52.40% on the 2-hop HOVER task, surpassing its CoT-only score of 50.02%. This pattern of improvement is consistent with GPT-3.5, where InfoRE and CoT together achieved the highest reported score of 70.24% on the same dataset. GPT-4 shows superior performance across all methods and datasets. With the Standard approach, it reaches 70.06% accuracy in the 4-hop challenge, suggesting an inherent advanced reasoning ability. This performance is mirrored in the specialized benchmarks, where GPT-4 with InfoRE attains near-perfect accuracy on FEVEROUS (94.67%) and very high accuracy on SCIFACT (93.55%). These figures solidify the notion that the InfoRE indeed enhances the reasoning capabilities of LLMs.
5.2 Question Answer
Different from claim verification datasets, this task involves using multiple documents as context, presenting a challenge in cross-document reasoning. Intuitively, when the reasoning process involves multiple documents, information re-organization can effectively merge the information from different documents and uncover logical relationships that are not apparent in plain text. Performance in Table 3 verifies the intuitive. We can see that after applying information re-organization, all the results have a significant improvement. GPT-4 outperforms the other models, with the highest F1 score of 76.27% when employing InfoRE on 2WikiMultiHopQA, and results on GPT-3.5 also show a notable increase in performance with InfoRE, achieving a 64.51% F1 score, surpassing its Standard and CoT results. The StrategyQA dataset, known for requiring strategic reasoning, GPT-4 with InfoRE reaching an F1 of 70.46%. These scores are the highest across all models and methods, with GPT3.5 also showing a solid increase with InfoRE at 62.78% F1. The HotpotQA results further affirm the superiority of GPT-4 using InfoRE, which achieves an F1 of 83.75%. This is consistent with the performance improvements observed with GPT3.5 (text-davinci-003) and Llama2-70B when applying InfoRE.
5.3 Machine Reading Comprehension
Different from conventional question-answering tasks, this task includes a set of candidate answers, and the goal of the model is to identify the correct answer from these options. This not only requires LLMs to deeply understand the contextual content but also to accurately match the answers with the available candidates. Moreover, understanding candidate answers is crucial because they are sometimes disruptive to correct answers, consequently increasing the difficulty of the reasoning task. Table 3 illustrates performance on the WIKIHOP and HotPotQA datasets. Our InfoRE consistently shows improvements across all datasets. Specifically, in the WIKIHOP dataset, GPT-3.5 with InfoRE outperforms other methods with an F1 of 51.13%. This improvement further verifies the effectiveness of our method.
6 Analysis
To verify that information re-organization enables LLMs to improve reasoning ability, we conduct a series of analysis experiments on the question answer and claim verification tasks. We first evaluate the quality of re-organized information with GPT-4 (gpt-4-32k) OpenAI (2023) in Section 6.1 Then we analyze the impact of the quality of re-organized information on reasoning in Section 6.2 and conduct a detailed error analysis of the predictions in Section 6.3.
6.1 Quality of Re-Organized Information
To assess the quality of the re-organized information, we perform a quantitative evaluation of the re-organized information with GPT-4 (gpt-4-32k) on the 2WikiMultiHopQA dataset. Specifically, we select 100 samples from the dataset, GPT-4 (gpt-4-32k) is asked to rank re-organized information produced by GPT-3.5 (text-davinci-003) Brown et al. (2020) and GPT-4, as well as original context following criteria: (1) Depth: The information present multiple relationships of a topic, offering insightful perspectives or in-depth understanding of the subject. (2) Clarity: Information is clear and precise, making it easy to understand without ambiguity, and contributes directly to the overall message. All of the information is ranked 1, 2, and 3 with 3, 2, and 1 scores, respectively. Finally, we get a weighted average score for each information to measure the overall quality.
Methods Depth
1st 2nd 3rd Avg R.
Original 0.22 0.36 0.42 1.80
GPT-3.5 0.34 0.35 0.31 2.03
GPT-4 0.44 0.29 0.27 2.17
Methods Clarity
1st 2nd 3rd Avg R.
Original 0.25 0.35 0.40 1.85
GPT-3.5 0.35 0.33 0.32 2.03
GPT-4 0.40 0.32 0.28 2.12
Table 4: Qualitative evaluation results on 2WikiMultiHopQA dataset. Avg R denotes the weighted average ranking score. The larger ranking score denotes better information quality.
Results are shown in Table 4. The results demonstrate that the re-organized information outperforms the original textual context in terms of depth and clarity, which justifies the motivation of our paper. In addition, the re-organized information from GPT-4 outperforms GPT-3.5, which proves in another way that GPT-4 is indeed more capable than GPT-3.5. The evaluation results further validate the effectiveness of our method. Furthermore, the 20.56% improvement in depth is more obvious than the 14.59% improvement in clarity, which indicates that the re-organization of the information has been particularly effective in enhancing the logical relationships of the information. The improvement in clarity suggests that the information is now presented in a more direct and streamlined manner, making it easier for LLMs to grasp the essential points without wading through unnecessary details.
6.2 Effect of Re-organized Information Quality
To explore the effects of the quality of re-organized context on model reasoning capabilities, we utilize both GPT-3.5 (text-davinci-003) and GPT-4 to re-organize context information. Reasoning processes are then independently executed on each model. Moreover, employing this cross-validation technique also allows us to effectively evaluate the robustness of our method.
Methods FEVEROUS 2WikiMultiHopQA
GPT-3.5 (†)
  Standard 92.00 58.25
  InfoRE 94.00 63.80
  InfoRE * 94.52 65.86
GPT-4 (*)
  Standard 92.33 72.69
  InfoRE 94.67 76.27
  InfoRE 94.32 74.82
Table 5: F1 performance of cross-validation, where InfoRE* denotes reason with GPT-3.5 but information re-organization with GPT-4, InfoRE denotes reason with GPT-4 but information re-organization with GPT-3.5 (text-davinci-003).
The cross-validation results, shown in Table 5, indicate that the use of either GPT-3.5 or GPT-4 for information re-organization leads to a marked improvement in the performance of LLMs on reasoning tasks. Additionally, it is observed that using GPT-4 for information re-organization while reasoning with the GPT-3.5 results in a further increase in reason results by 0.52% and 2.06% on FEVEROUS and 2WikiMultiHopQA datasets, respectively. This implies that enhancing the quality of re-organized information can improve the performance of LLMs, pointing to a promising direction for future research in refining information synthesis and re-organization. Conversely, when GPT-3.5 is employed for information re-organization in conjunction with reasoning using the GPT-4 model, there is a decrease in reason ability by 0.35% and 1.45% respectively. In addition, the magnitude of the decrease is lower than the increase, which implies that our re-organization strategy may have a greater impact on models with weaker reasoning abilities. This finding aligns with our understanding that models with inherently weaker reasoning abilities tend to rely more heavily on external strategy. In contrast, models with stronger inherent capabilities can use their strengths to mitigate the adverse effects.
6.3 Error Analysis
To better comprehend where the errors in our InfoRE methodology come from and where they are fixed, we annotate 100 wrong predictions made by both InfoRE and Standard methods with GPT-3.5 (text-davinci-003) on 2WikiMultiHopQA dataset. We categorize the errors into $4$ classes:
  • Contextual Misunderstanding This happens when the model fails to interpret or connect multiple pieces of information from different parts of the documents. Multi-hop reasoning requires synthesizing information from various segments, and recognizing logical relations, and any misunderstanding can lead to incorrect conclusion.
  • Factual Error The model may provide an answer that is factually incorrect or not supported by the given documents. This is often due to the model’s reliance on its training data, which may not always align with the specific facts in the context.
  • Mathematical Error The error occurs when math calculations are involved in deriving the final answer.
  • Unanswerable Question It’s a specific type of error or limitation in dataset design, where the context does not contain enough information to provide a valid answer to the posed question.
When engaging in reasoning with large language models, all four error categories are present. As shown in Figure 3, there are 6% unanswerable question errors in the dataset, and more than 90% errors occur in the reasoning in the baseline method. Among the four types of error, contextual misunderstanding is the primary source of errors in the baseline, highlighting the importance of an in-depth understanding of context in solving reasoning tasks with large language models. This finding is consistent with our motivation presented in the introduction section. Moreover, comparing the results of errors between our method and the baseline method, our method mainly corrects 13% of errors coming from the baseline method, most of the corrected errors are contextual misunderstanding errors. This further indicates that our InfoRE method assists Large Language Models (LLMs) in understanding context, signifying the necessity and effectiveness of conducting information re-organization before directly addressing the original question.
Figure 3: Error Analysis of InfoRE on 2WikiMultiHopQA against Standard baseline method. The first four rectangles are error categories, while “Corrected” on the far right denotes the percentage of errors originally made by the baseline method that our method InfoRE has successfully corrected.
7 Conclusion
In this paper, we propose an information re-organization method to improve the reasoning ability of large language models (LLMs). Compared with previous approaches primarily focus on improving the quality of intermediate steps, our method emphasizes uncovering the logical relationships and multi-hop connections by information re-organization. By re-organizing the way information is presented, we aim to enable LLMs to explicitly perceive the logical relationships and multi-hop connection of concepts in the given context. This facilitates a deeper integration and understanding of context, leading to more robust reasoning results. To verify the effectiveness of our method, we conduct experiments using various LLMs across a range of contextually aware multi-hop reasoning tasks. The experiment results demonstrate the potential of our method to improve the reasoning ability of LLMs.
Appendix A Prompt for Obtaining MindMap
The specific prompt to obtain MindMap is shown in Table  6. In the prompt, the content within the brackets should be replaced with specific example content.
Prompt Content for Obtaining MindMap
Given a claim and corresponding evidence, please summarize the evidence as a mind map according to the claim. The output must be in a strict JSON format: {“mind_map”: “mind_map”}. CLAIM: [CLAIM] EVIDENCE: [EVIDENCE] .
Table 6: Specific prompt for obtaining MindMap. During execution, “[EVIDENCE]” needs to be replaced with the specific document, and “[CLAIM]” is replaced with the specific question.
Appendix B Prompt for Multi-hop Reason
The specific prompts for Standard, chain-of-thought (CoT), our InfoRE, and InfoRE + CoT to reason are shown in Table  7. In the prompt, the content within the brackets [] should be replaced with specific example content. Specifically, in the claim verification task, [CLAIM] corresponds to the claim of the specific example, and [EVIDENCE] corresponds to the evidence of the specific example. In the question-answering task, [CLAIM] corresponds to the question of the specific example, and [EVIDENCE] corresponds to the context document. [MindMap] corresponds to the MindMap of the specific example.
Methods Prompt Content
Standard Documents: [EVIDENCE] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Answer:
CoT Documents: [EVIDENCE] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Let’s think step by step. Answer:
InfoRE Documents: [MindMap] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Answer:
InfoRE + CoT Documents: [MindMap] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Let’s think step by step. Answer:
Table 7: Specific prompts and answer format instructions of Standard, CoT, InforRE, and InfoRE + CoT in our paper. During execution, “[EVIDENCE]” needs to be replaced with the specific document, “[CLAIM]” is replaced with the specific question, and “[MindMap]” is replaced with the specific MindMap.