### v2

# Information Re-Organization Improves Reasoning in Large Language Models

Abstract
Improving the reasoning capabilities of large language models (LLMs) has attracted considerable interest. Recent approaches primarily focus on improving the reasoning process to yield a more precise final answer. However, in scenarios involving contextually aware reasoning, these methods neglect the importance of first identifying logical relationships from the context before proceeding with the reasoning. This oversight could lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning outcomes. In this paper, we propose an information re-organization (InfoRE) method before proceeding with the reasoning to enhance the reasoning ability of LLMs. Our re-organization method involves initially extracting logical relationships from the contextual content, such as documents or paragraphs, and subsequently pruning redundant content to minimize noise. Then, we utilize the re-organized information in the reasoning process. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships, while also ensuring high-quality responses by eliminating potential noise. To demonstrate the effectiveness of our approach in improving the reasoning ability, we conduct experiments using Llama2-70B, GPT-3.5, and GPT-4 on various contextually aware multi-hop reasoning tasks. Using only a zero-shot setting, our method achieves an average absolute improvement of 4% across all tasks, highlighting its potential to improve the reasoning performance of LLMs ***https://github.com/hustcxx/InfoRE.
22footnotetext:   Corresponding author.
1 Introduction
Large language models (LLMs) demonstrate powerful generative capabilities and achieve remarkable performance across a range of linguistic tasks [1, 2, 3]. However, their capabilities in performing complex reasoning tasks — an essential aspect of advanced language understanding and intelligent decision-making — still present substantial challenges [4, 5]. This has spurred researchers to explore innovative strategies [6, 7, 8] to improve the reasoning capabilities of these models.
Recently, diverse methods have been developed to enhance the reasoning ability of LLMs. For example, a notable method Chain-of-Thought (CoT) [6], incorporates a series of intermediate reasoning steps into the reasoning. CoT [6] allows for a more transparent and understandable path to the final answer, making it easier to follow the logic behind the conclusion. Building upon this foundation, subsequent approaches such as Tree of Thoughts (ToT) [8] and Graph of Thoughts (GoT) [8] are proposed to further refine the reasoning steps and enhance the accuracy and reliability of LLMs. Different from the sequential intermediate steps of Chain-of-Thought [6], Tree of Thoughts [8] and Graph of Thoughts [8] model the problem solving process into structured tools of tree and graph, respectively. A critical observation is that these existing approaches primarily focus on improving the reasoning process of large language models, as shown in Figure 1 (left). However, in scenarios involving contextually aware reasoning, it is equally important to first identify logical relationships of context before proceeding with the reasoning, not just improve their reasoning process. This is because logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of reasoning [9]. Nevertheless, these existing methods often neglect this crucial step. Such an oversight can lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning results.
Figure 1: InfoRE (Ours) vs existing methods. In contrast to the existing methods that primarily focus on the reasoning process, our InfoRE emphasizes the re-organization of context information.
Inspired by the fact that when faced with context-aware reasoning tasks humans often first re-organize existing contextual information to uncover the logical relationships, eliminate noises, and enhance their understanding of the context, we propose an information re-organization (InfoRE) method, to ground reasoning by the re-organized information. As shown in Figure 1 (right), different from previous methodologies that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, our approach takes a novel direction of context re-organization. We emphasize the utilization of re-organized contextual content to explicitly present the logical relationships that are often implicit within the plain text, promoting more effective reasoning. Specifically, our re-organization method comprises two operations: extraction and pruning. The extraction first uncovers the implicit logical relationships within the contextual content by transforming the content into a MindMap structure [10]. We employ this structure because it is rich in logical relationships and encompasses multi-hop connections. Pruning is then used to further minimize noise that is irrelevant to the reasoning objective. The pruning operation uses a pre-trained BERT [11] based model trained with reinforcement learning (RL). Finally, we utilize the re-organized context to reason. This enables LLMs to deeply understand the context by clearly perceiving these logical relationships, facilitating the quality and reliability of reasoning. Besides, our information re-organization method can be integrated with existing prompt methods, like CoT [6], to further improve the reasoning ability of LLMs. To verify the efficacy of our proposed InfoRE method, we conduct experiments using Llama2-70B [2], GPT-3.5 [1], and GPT-4 [3] on various contextually aware multi-hop reasoning tasks, including claim verification [12], question answering [13], and reading comprehension [14]. Using only a zero-shot setting, our method achieves an average improvement of 4% across all tasks, highlighting its potential to improve the reasoning performance of LLMs.
Our main contributions are as follows:
  • In contrast to existing methods that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, we take a novel direction of context re-organization.
  • The re-organization method initially uncovers the logical relationships that encompass multi-hop connections in the contextual content by extraction, and subsequently minimizes noise by pruning.
  • Experiment improvements on contextually aware multi-hop reasoning tasks across claim verification, question answering, and reading comprehension show the efficacy of our proposed method.
2 Related Work
Reasoning with LLMs
Large language models (LLMs) [2, 1, 3], have revolutionized the field of natural language processing (NLP) and demonstrate remarkable proficiency across a range of linguistic tasks. To further improve the reasoning ability of LLMs has attracted considerable interest. In-context learning [1], as a promising approach to enhance the reasoning abilities of LLMs, has been verified in mathematical reasoning task [6]. In addition, Chain of Thought (CoT) [6] incorporates a coherent series of intermediate reasoning steps to improve the reasoning ability of LLMs. Following this paradigm, Tree-of-Thoughts (ToT) [7] and Graph-of-Thoughts (GoT) [8] have also been proposed to focus on improving the structure of the intermediate chain. Then self-consistency [15] and plan-to-solve [15] focus on the reliability of the chain. Recently, the step-back prompting [16] is proposed, which obtains the high-level concept and first principles from instances by abstraction in the first step, then guides the reasoning of LLMs with the obtained concept and principles. For tasks that require multi-hop reasoning, current approaches [17, 18] generally decompose multi-hop problems into simpler sub-tasks problems relying on the in-context learning ability of LLMs. In contrast to existing methods, we take a novel direction of context re-organization to enhance the reasoning capabilities of LLMs for contextually aware reasoning tasks.
Information Re-organization
Information re-organization is a technique that leverages other structures to improve the clarity and comprehensibility of information. Moreover, information re-organization can also reveal conceptual relationships implicit in the original textual text. The graph, as the most common re-organization structure, is often used to enhance contextual representation in a variety of tasks. For example, BASS [19] re-organizes the original document to a semantic graph to improve the summary generation. DGM [20] re-organizes the context to an explicit discourse graph in the reading comprehension task. MindMap [10] is a powerful structure method for representing knowledge and concepts, it can be used to construct a hierarchical abstraction of natural language text [21]. Previous method [22] uses it to organize a large amount of scientific material to enhance the clarity of the text material. In this paper, we use it as the structure of re-organized information to uncover the logical relationships and multi-hop connections implicit within the plain context. Then we use this re-organized information to further improve the reasoning ability of LLMs.
3 Methodology
In this section, we first give a formulation of a context-aware reasoning task in §3.1 and then describe our method in detail. As shown in Figure 2, the framework of our method consists of two components, an information re-organization §3.2 and a reasoning step using the re-organized context §3.3.
Figure 2: Illustration of our information re-organization method with two modules: 1)Information Re-Organization, which includes logic relationship extraction and noise pruning. 2) Reasoning using re-organized context. The re-organized context in black italicized text is relevant to the question.
3.1 Task Formulation
Given a sample $(c,q,a)$ from sample space $(\mathcal{C},\mathcal{Q},\mathcal{A})$, where $c$ denotes gold context, $q$ denotes question, the task aims to obtain answer $a$ to the question $q$, based on gold context $c$ with a large language model. In particular, it requires multi-hop reasoning to get the answer $a$ to question $q$.
3.2 Information Re-Organization
The purpose of the information re-organization is to obtain the logical relationships from the context and minimize the noise irrelevant to the question. This goal is achieved through two operations: extraction and pruning.
3.2.1 Extraction
For a question $q$, if the derivation of its answer $a$ relies on context $c$, then a deeper understanding of $c$ is crucial. Logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of understanding and reasoning [9]. However, the logical relationships in the plain context are often implicit. Therefore, we perform a extraction operation on the plain context to uncover the logical relations in it using a language model. The process can be defined as:
$g=f_{\theta}(c,q,P_{in})$ (1)
where $f_{\theta}(\cdot)$ is a language model parameterized by $\theta$, $P_{in}$ is input task prompt. The specific task prompt $P_{in}$ used in our paper is displayed in the Appendix A. We perform Equation 1 for each $c$ in the sample space to obtain the extracted context space $\mathcal{G}$. We use the MindMap [10] structure to display the reorganized content, because it serves as a powerful structure for representing knowledge, concepts, and perspectives, contains not only logical relationships but also multi-hop connections.
[…]
3.2.2 Pruning
As described in Section 3.2, the extracted context $g\in\mathcal{G}$ contains various logical relationships and rich attributes. However, not all logical relationships and attributes help answer the question $q$. On the contrary, some may even interfere with the response to the question. For example, consider the question in Figure 2, "Where did the producer of Julius Caesar study or work?". In this case, the content "Julius Caesar $\rightarrow$ Production Company" is a distracting element, and "Julius Caesar $\rightarrow$ Adaptation" is irrelevant to the question.
Figure 3: Illustration of Pruning model. The representation of [CLS] is used to obtain action probabilities.
To further reduce the interference of distracting or irrelevant logical relations and attributes on the retrieval of answers for question $q$, we use a pruning model trained through reinforcement learning (RL). The pruning model is based on the pre-trained BERT [11], as shown in Figure 3. Its input consists of concatenated logical relationships, their corresponding attribute values from $g$, and the question $q$. For example, to prune the relation Julius Caesar $\rightarrow$ Adaptation, the input is: [CLS] Julius Caesar Adaptation Play by Shakespeare [SEP] Question [SEP]. We give a detailed demonstration of input format in Appendix B.
RL Formulation
We formulate the pruning policy model optimization as an RL problem and employ proximal policy optimization (PPO) [23]. The action is keeping or deleting the logical relations. The policy decides the action probability given the question and $g$. We fine-tune the policy model $\pi$ by optimizing the reward $r$:
$\mathbb{E}_{\pi}[r]=\mathbb{E}_{\boldsymbol{g}\sim\mathcal{G},\boldsymbol{q}\sim\mathcal{Q},\boldsymbol{z}\sim\pi(\cdot\mid\boldsymbol{x},\boldsymbol{q})}[r(\boldsymbol{z},\boldsymbol{q})]$ (2)
Reward Function
Our goal is to maximize LLM’s generation toward the desired target by an alignment measure $\mathcal{R}$, and we use this as a reward. In this paper, the alignment metric we have chosen is the F1 score. To keep the policy network $\pi$ from moving too far from the old result, we add a clipped term in the reward. Therefore, the final reward becomes:
$r(\boldsymbol{z},\boldsymbol{q})=\mathrm{min}(\mathcal{R}(\boldsymbol{z},\boldsymbol{q}),\mathrm{clip}(\pi(\boldsymbol{z}\mid\boldsymbol{x},\boldsymbol{q}),1-\epsilon,1+\epsilon))$ (3)
where $\epsilon$ is a hyperparameter indicating the range of CLIP to be performed. After pruning, the extracted contextual content $g$ becomes the context $g^{{}^{\prime}}$, which is closely related to the reasoning objective $q$. We perform a pruning operation for each $g\in\mathcal{G}$ to obtain the pruned context space $\mathcal{G}^{{}^{\prime}}$.
3.3 Reasoning
After the re-organization process in Section 3.2, we get the re-organized context $g^{{}^{\prime}}\in\mathcal{G}^{{}^{\prime}}$. Then the re-organized context $g^{{}^{\prime}}$ can be used as a context alone or combined with the original context $c$ to get the final answer of $q$. The reasoning process can be defined as:
$o=f_{\theta}(g^{{}^{\prime}},\left[c\right],q,P_{r})$ (4)
[…]
4 Experimental
4.1 Tasks and Datasets
To verify the effectiveness of our information re-organization method, we conduct experiments across a range of contextually aware multi-hop reasoning tasks and datasets, including claim verification [12], question answering [13], and reading comprehension [14]. The detailed dataset information, including data splits and statistics, is available in Appendix D.
Claim Verification The task involves assessing a given claim against a set of evidence documents to determine whether they support or refute the claim [12]. We consider HOVER [24] and FEVEROUS [25], which comprise complex claims that necessitate multi-hop reasoning for verification. Besides, we also take into account the SCIFACT [26] dataset, notable for its inclusion of scientific claims.
Question Answering For this task, we consider the following datasets: 2WikiMultiHopQA [27], StrategyQA [28], MuSiQue [29], and HotpotQA [30]. To answer the questions in these datasets requires not only multi-hop reasoning but also cross-document analysis.
Reading Comprehension Machine reading comprehension task requires a model to process documents and select an answer from the provided candidates to a question about the content [14]. We primarily consider WIKIHOP [31] in the task, which necessitates multi-hop reasoning to derive the final answer. Additionally, HotpotQA [30] is also frequently considered as part of the QA domain.
4.2 Baselines
In our paper, we compare our method InfoRE to two reasoning baselines: 1) Standard. 2) CoT. The Standard approach is a method that directly reasons with the original textual context. The Chain of Thought (CoT) [6] method involves augmenting standard reasoning methods by adding a step-by-step thought process. In our paper, we conduct the CoT strategy by appending the sentence “Let’s think step by step.” at the end of the question.
The baseline methods and our InfoRE both adopt a zero-shot setting to counteract the potential randomness associated with demonstrations in a few-shot setting. We also design an answer-format instruction within the prompts for various tasks to standardize the structure of the final answer, thereby enhancing the precision of answer extraction. Moreover, all results reported in the paper use only the reorganized contextual information to reason. Comprehensive details about prompts and answer-format instruction are available in Appendix C. Following previous methods [17, 18, 25], we run the official evaluation scripts of each dataset to get the F1 to measure the results.
4.3 Implementation Details
In our paper, the large language models employed in the extraction and reasoning process include Llama2-70B [2], GPT-3.5 (text-davinci-003) [32] and GPT-4 [3]. We configure all models with top_p parameter as $1.0$ and temperature as $0.0$. In the policy model, we use the BERT-base version on all tasks and datasets. In RL training, we calculate the F1 score between the generated answer and the reference answer as the reward, with a rescaling coefficient of 10. We train the model for 1000 episodes. We conduct training for epoch 5, a batch size of 4, and a learning rate of 2e-6. The parameter of $\epsilon$ is set to 0.2. All experiments are conducted on an NVIDIA RTX A6000.
5 Results and Analysis
5.1 Main Results
Claim Verification
Table 1 presents a comprehensive performance comparison between our InfoRE and existing zero-shot techniques. For the HOVER dataset, we segment it into the 2-hop, 3-hop, and 4-hop levels following previous methods [17]. As depicted in Table 1, our InfoRE demonstrates significant improvements in the zero-shot claim verification task. The CoT [6] approach offers a lightly increase of 0.62% in the HOVER 4-hop using GPT-4, which indicates its marginal utility in more complex reasoning scenarios. Yet, our InfoRE achieves 3.02% improvement on the HOVER 4-hop using GPT-4 showing remarkable performance on contextual aware understanding and reasoning. This improvement is further increased to 73.62% in combination with CoT, suggesting that the methods complement each other effectively. In the case of Llama2-70B, the combined application of InfoRE and CoT yields a score of 53.20% on the 2-hop HOVER task, surpassing its CoT-only score of 50.02%. This pattern of improvement is consistent with GPT-3.5. GPT-4 shows superior performance across all methods and datasets, suggesting an inherent advanced reasoning ability. This performance is mirrored in the specialized benchmarks, where GPT-4 with InfoRE attains near-perfect accuracy on FEVEROUS (95.62%) and very high accuracy on SCIFACT (93.67%). These figures solidify the notion that the InfoRE indeed enhances the reasoning capabilities of LLMs.
Table 1: Zero-shot performance on claim verification task across three large language models.
LLMs Methods HOVER FEVEROUS SCIFACT
2-hop 3-hop 4-hop
LLAMA2-70B Standard 49.41 48.35 47.82 63.39 60.70
InfoRE 52.83 51.42 50.04 67.84 63.81
$\uparrow 3.42$ $\uparrow 3.07$ $\uparrow 2.22$ $\uparrow 4.45$ $\uparrow 3.11$
CoT 50.02 48.76 48.01 64.53 61.24
InfoRE + CoT 53.20 51.70 50.15 68.12 64.02
$\uparrow 3.18$ $\uparrow 2.94$ $\uparrow 2.14$ $\uparrow 3.59$ $\uparrow 2.78$
GPT-3.5 Standard 64.74 63.04 61.54 87.67 77.42
InfoRE 68.21 66.45 64.91 91.31 81.54
$\uparrow 3.47$ $\uparrow 3.41$ $\uparrow 3.37$ $\uparrow 3.64$ $\uparrow 4.12$
CoT 66.70 64.52 62.69 88.67 78.49
InfoRE + CoT 69.02 67.53 65.66 91.53 82.26
$\uparrow 2.32$ $\uparrow 3.01$ $\uparrow 2.97$ $\uparrow 2.86$ $\uparrow 3.77$
GPT-4 Standard 72.40 71.02 70.06 92.33 91.40
InfoRE 75.87 74.06 73.08 95.62 93.67
$\uparrow 3.47$ $\uparrow 3.04$ $\uparrow 3.02$ $\uparrow 3.29$ $\uparrow 2.27$
CoT 73.82 72.07 70.68 92.67 92.47
InfoRE + CoT 76.69 75.16 73.62 95.67 94.32
$\uparrow 2.87$ $\uparrow 3.09$ $\uparrow 2.94$ $\uparrow 3.00$ $\uparrow 1.85$
Question Answer and Reading Comprehension
Different from claim verification task, this task involves using multiple documents as context, presenting a challenge in cross-document reasoning. Intuitively, when the reasoning process involves multiple documents, information re-organization can effectively merge the information from different documents and uncover logical relationships that are not apparent in plain text. Performance in Table 2 verifies the intuitive. We can see that after applying information re-organization, all the results have a significant improvement. GPT-4 outperforms the other models, with the highest F1 score of 76.52%, 71.20%, and 83.22% when employing InfoRE on 2WikiMultiHopQA, StrategyQA, and HotpotQA, respectively. This is consistent with the performance improvements observed with GPT3.5 (text-davinci-003) and Llama2-70B when applying InfoRE. Different from conventional QA tasks, reading comprehension tasks require LLMs to not only deeply understand the context but also identify distractors among the candidates, increasing the reasoning challenge. Our InfoRE consistently shows improvements on this task. Specifically, in the WIKIHOP dataset, GPT-3.5 with InfoRE outperforms other methods with an F1 of 51.87%. This improvement further verifies the effectiveness of our method.
Table 2: Zero-shot results on Question Answering and Reading Comprehension tasks. 2WMHQA, SQA, and HQA are abbreviations for 2WikiMultiHopQA, StrategyQA, and HotpotQA, respectively.
LLMs Methods 2WMHQA MuSiQue SQA HQA WIKIHOP
LLAMA2 (70B) Standard 52.56 49.55 51.23 66.07 40.32
InfoRE 57.62 52.78 55.32 69.98 42.90
$\uparrow 5.06$ $\uparrow 3.23$ $\uparrow 4.09$ $\uparrow 3.91$ $\uparrow 2.58$
CoT 52.99 52.90 56.80 66.80 41.07
InfoRE + CoT 57.72 56.10 59.93 70.60 43.37
$\uparrow 4.73$ $\uparrow 3.20$ $\uparrow 3.13$ $\uparrow 3.80$ $\uparrow 2.30$
GPT-3.5 Standard 58.25 55.01 59.39 73.30 48.92
InfoRE 64.58 58.03 63.16 77.12 51.87
$\uparrow 6.33$ $\uparrow 3.02$ $\uparrow 3.77$ $\uparrow 3.82$ $\uparrow 2.95$
CoT 59.37 57.05 67.51 73.90 49.65
InfoRE + CoT 65.13 60.52 70.45 77.74 52.70
$\uparrow 5.76$ $\uparrow 3.47$ $\uparrow 2.94$ $\uparrow 3.84$ $\uparrow 3.05$
GPT-4 Standard 72.69 62.65 68.32 79.33 55.46
InfoRE 76.52 66.36 71.20 83.22 58.01
$\uparrow 3.83$ $\uparrow 3.71$ $\uparrow 2.88$ $\uparrow 3.89$ $\uparrow 2.55$
CoT 74.08 64.36 68.50 80.66 56.02
InfoRE + CoT 78.60 69.11 71.54 84.26 58.91
$\uparrow 4.52$ $\uparrow 4.75$ $\uparrow 3.04$ $\uparrow 3.60$ $\uparrow 2.89$
5.2 Analysis
Ablation studies
In our paper, the re-organization comprises two components: extraction and pruning. To investigate the impact of each component in detail, we conduct a series of ablation experiments using GPT-3.5 on the 2WikiMultiHopQA dataset. First, we directly remove extraction and pruning from our method, and the results are shown in the second and third rows of Table 3, respectively. It is worth mentioning that in the experiment where extraction is removed, we directly prune the sentences in the original context. Furthermore, we replace the reinforcement learning-based pruning method with a similarity-based pruning method to demonstrate its effectiveness. Specifically, the similarity-based pruning method uses Siamese-BERT, which takes the original question and each logical relationship as inputs separately, and then generates the corresponding representations. Then, we calculate the cosine similarity between these two representations. Finally, we removed the 30% of logical relationships with the lowest similarity, the result is shown in the last row of Table 3.
Table 3: F1 performance of ablation studies.
Methods 2WikiMultiHopQA
Full model 64.58
w/o extraction 61.64
w/o pruning 63.05
similarity-based pruning 63.32
The results in Table 3 show that removing the extraction and pruning operations leads to performance drops of 2.94% and 1.53%, respectively. This demonstrates the effectiveness of both components in our methods. The larger performance drop after removing extraction highlights the importance of extracting logical relationships for effective reasoning. After replacing the pruning model, the performance dropped by 1.26%, but the results were still better than without pruning. This not only demonstrates the necessity of pruning but also highlights the effectiveness of the reinforcement learning-based pruning method.
Quality of Re-Organized Information
To assess the quality of the re-organized information, we perform a quantitative evaluation of the re-organized information with GPT-4 (gpt-4-32k) on the 2WikiMultiHopQA dataset. Specifically, we select 100 samples from the dataset, GPT-4 (gpt-4-32k) is asked to rank re-organized information produced by GPT-3.5 (text-davinci-003) [1] and GPT-4, as well as original context following criteria: (1) Depth: The information present multiple relationships of a topic, offering insightful perspectives or in-depth understanding of the subject. (2) Clarity: Information is clear and precise, making it easy to understand without ambiguity. All of the information is ranked 1, 2, and 3 with 3, 2, and 1 scores, respectively. Finally, we get a weighted average score for each information to measure the overall quality.
Table 4: Qualitative evaluation results on 2WikiMultiHopQA dataset. Avg R denotes the weighted average ranking score. The larger ranking score denotes better information quality.
Methods Depth
1st 2nd 3rd Avg R.
Original 0.22 0.36 0.42 1.80
GPT-3.5 0.32 0.36 0.32 2.00
GPT-4 0.46 0.28 0.26 2.20
Methods Clarity
1st 2nd 3rd Avg R.
Original 0.25 0.35 0.40 1.85
GPT-3.5 0.35 0.32 0.33 2.02
GPT-4 0.40 0.33 0.27 2.13
The results in Table 4 demonstrate that the re-organized information outperforms the original textual context in terms of depth and clarity, which justifies the motivation of our paper. In addition, the re-organized information from GPT-4 outperforms GPT-3.5, which proves in another way that GPT-4 is indeed more capable than GPT-3.5. The evaluation results further validate the effectiveness of our method. Furthermore, the 22.22% improvement in depth is more obvious than the 15.14% improvement in clarity, which indicates that the re-organization of the information has been particularly effective in enhancing the logical relationships of the information. The improvement in clarity suggests that the information is now presented in a more direct and streamlined manner, making it easier for LLMs to grasp the essential points without wading through unnecessary details.
Effect of Re-organized Information Quality
To explore the effects of the quality of re-organized context on model reasoning capabilities, we utilize both GPT-3.5 (text-davinci-003) and GPT-4 to re-organize context information. Reasoning processes are then independently executed on each model. Moreover, employing this cross-validation technique also allows us to effectively evaluate the robustness of our method.
Table 5: F1 performance of cross-validation, where InfoRE* denotes reason with GPT-3.5 but information re-organization with GPT-4, InfoRE denotes reason with GPT-4 but information re-organization with GPT-3.5 (text-davinci-003).
Methods FEVEROUS 2WikiMultiHopQA
GPT-3.5 (†)
    Standard 87.67 58.25
    InfoRE 91.31 64.58
    InfoRE * 92.50 66.61
GPT-4 (*)
    Standard 92.33 72.69
    InfoRE 95.62 76.52
    InfoRE 94.67 75.07
The cross-validation results, shown in Table 5, indicate that the use of either GPT-3.5 or GPT-4 for information re-organization leads to a marked improvement in the performance of LLMs on reasoning tasks. Additionally, it is observed that using GPT-4 for information re-organization while reasoning with the GPT-3.5 results in a further increase in reason results by 1.19% and 2.03% on FEVEROUS and 2WikiMultiHopQA datasets, respectively. This implies that enhancing the quality of re-organized information can improve the performance of LLMs, pointing to a promising direction for future research in refining information synthesis and re-organization. Conversely, when GPT-3.5 is employed for information re-organization in conjunction with reasoning using the GPT-4 model, there is a decrease in reason ability by 0.95% and 1.45% respectively. In addition, the magnitude of the decrease is lower than the increase, which implies that our re-organization strategy may have a greater impact on models with weaker reasoning abilities. This finding aligns with our understanding that models with inherently weaker reasoning abilities tend to rely more heavily on external strategy. For models with stronger inherent capabilities, our method still further improve its reasoning ability.
Error Analysis
To better comprehend where the errors in our InfoRE methodology come from and where they are fixed, we annotate 100 wrong predictions made by both InfoRE and Standard methods with GPT-3.5 on 2WikiMultiHopQA dataset. We categorize the errors into $4$ classes:
  • Contextual Misunderstanding (CM): This happens when the model fails to interpret or connect multiple pieces of information from different parts of the documents. Multi-hop reasoning requires synthesizing information from various segments, and recognizing logical relations, and any misunderstanding can lead to incorrect conclusion.
  • Factual Error (FE): The model may provide an answer that is factually incorrect or not supported by the given documents. This is often due to the model’s reliance on its training data, which may not always align with the specific facts in the context.
  • Mathematical Error (ME): The error occurs when math calculations are involved in deriving the final answer.
  • Unanswerable Question (UQ): It’s a specific type of error or limitation in dataset design, where the context does not contain enough information to provide a valid answer to the posed question.
Figure 4: Error Analysis of InfoRE on 2WikiMultiHopQA against Standard baseline method. The first four rectangles are error categories, while “Corrected" on the far right denotes the percentage of errors originally made by the baseline method that our method InfoRE has successfully corrected.
When engaging in reasoning with large language models, all four error categories are present. As shown in Figure 4, there are 6% unanswerable question errors in the dataset, and more than 90% errors occur in the reasoning in the baseline method. Among the four types of error, contextual misunderstanding is the primary source of errors in the baseline, highlighting the importance of an in-depth understanding of context in solving reasoning tasks with large language models. This finding is consistent with our motivation presented in the introduction section. Moreover, comparing the results of errors between our method and the baseline method, our method mainly corrects 14% of errors coming from the baseline method, most of the corrected errors are contextual misunderstanding errors. This further indicates that our InfoRE method assists Large Language Models (LLMs) in understanding context, signifying the necessity and effectiveness of conducting information re-organization before directly addressing the original question.
6 Conclusion
In this paper, we propose an information re-organization method to improve the reasoning ability of large language models (LLMs). Compared with previous approaches primarily focus on improving the quality of intermediate steps, our method emphasizes uncovering the logical relationships, multi-hop connections, and pruning the irrelevant information through information re-organization. This approach enables LLMs to explicitly perceive the logical relationships and multi-hop connections of concepts within the context, promoting a deeper integration and understanding of the context, which results in more robust reasoning outcomes. To verify the effectiveness of our method, we conduct experiments using various LLMs across a range of contextually aware multi-hop reasoning tasks. The experiment results demonstrate the potential of our method to improve the reasoning ability of LLMs. Additionally, our method has a positive impact on various tasks involving context understanding, such as academic research, legal analysis, and medical diagnostics. However, it is also important to be aware of potential negative impacts, such as the propagation of misinformation.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A Prompt for Extraction
The specific prompt of extraction is shown in Table  6. The purpose of the prompt is to obtain logical relationships and multi-hop connections from the context. In the prompt, the content within the brackets should be replaced with specific example content.
Table 6: Specific prompt for obtaining logical relationships. During execution, “[EVIDENCE]" needs to be replaced with the specific document, and “[CLAIM]" is replaced with the specific question.
Prompt Content for Logical Relationship Extraction
Given a claim and corresponding evidence, please summarize the evidence as a mind map according to the claim. The output must be in a strict JSON format: {“mind_map": “mind_map"}. CLAIM: [CLAIM] EVIDENCE: [EVIDENCE] .
Appendix B Input Format For Pruning
The extracted context $g$ includes logical relationships and corresponding attribute values. First, we iterate through all logical relationships and attribute values, and following the previous method [33], we concatenate them using [SEP] [CLS], combining them with the question to input into the pruning model. Then, we use the representation of the [CLS] token to represent the logical relationships for subsequent operations.
We use an example in Figure 2 to demonstrate the input format of the extracted context $g$ in the pruning model. The question is: Where did the producer of Julius Caesar study or work? The the extracted context $g$ is:
Julius Caesar:
  Production Company: Metro-Goldwyn-Mayer
  Director: Joseph L. Mankiewicz
  Producer:
     Name: John Houseman
     Education: Clifton College, England
     Occupation: Grain Trade, London
  Adaptation: Play by Shakespeare
We only traverse logical relationships of the first-level progressive type. Therefore, the logical relationships contained in extracted context $g$ of example include 1) Production Company: Metro-Goldwyn-Mayer, 2) Director: Joseph L. Mankiewicz, 3) Producer: Name: John Houseman, 4) Producer: Education: Clifton College,England, 5) Producer: Occupation: Grain Trade, London, 6) Adaptation: Play by Shakespeare. Then, we concatenate them using [SEP][CLS].
The input format after concatenation is: [CLS] Production Company: Metro-Goldwyn-Mayer [SEP] [CLS] Director: Joseph L. Mankiewicz [SEP] [CLS] [CLS] Producer: Name: John Houseman [SEP] [SEP] [CLS] Producer: Education: Clifton College [SEP] [CLS] Producer: Occupation: Grain Trade, London [SEP] [CLS] Adaptation: Play by Shakespeare [SEP] [CLS] Where did the producer of Julius Caesar study or work? [SEP].
Appendix C Prompt for Multi-hop Reason
During the reasoning stage using large language models, to accommodate more context-aware reasoning tasks while ensuring comparability of results, we designed a universal prompt template. The prompt template consists of three components: original context, e.g., documents or paragraphs, reorganized information, and a question. The prompt template is as follows:
Documents:
   [Re-Organized TEXT]
   [TEXT][Optional]
Question:
   [QUESTION]
please answer the question based on the documents.
Answer:
The specific prompts for Standard, chain-of-thought (CoT), our InfoRE, and InfoRE + CoT to reason are shown in Table  7. In the prompt, the content within the brackets [] should be replaced with specific example content.
Table 7: Specific prompts and answer format instructions of Standard, CoT, InforRE, and InfoRE + CoT in our paper. During execution, “[EVIDENCE]" needs to be replaced with the specific document, “[CLAIM]" is replaced with the specific question, and “[MindMap]" is replaced with the specific MindMap.
Methods Prompt Content
Standard Documents: [EVIDENCE] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Answer:
CoT Documents: [EVIDENCE] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Let’s think step by step. Answer:
InfoRE Documents: [MindMap] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Answer:
InfoRE + CoT Documents: [MindMap] Question: [CLAIM]? Please answer the question based on Documents. Your final answer should be enclosed in XML tag <answer></answer>, like this: <answer>{final_answer}</answer>, at the end of your response. Let’s think step by step. Answer:
Appendix D Detailed Dataset Information
In our experiments, due to the resource limitations of large language models, we sample a portion from each dataset, following previous methods [6, 16].
Table 8: Details statics information of evaluation datasets we used in the paper. Pairs denote the number of examples. 2WMHopQA is the short name of 2WikiMultiHopQA.
Tasks Dataset Pairs
train test
Claim Verification FEVEROUS 2000 2959
HOVER 2000 4000
SCIFACT 200 212
Question Answering 2WikiMultiHopQA 2000 500
StrategyQA 1000 229
MusiQue 2000 2417
Reading Comprehension HotpotQA 2000 500
WIKIHOP 2000 500
Claim Verification
The task involves assessing a given claim against a set of evidence documents to determine whether they support or refute the claim [12]. We consider HOVER [24] and FEVEROUS [25], which comprise complex claims that necessitate multi-hop reasoning for verification. Besides, we also take into account the SCIFACT [26] dataset, notable for its inclusion of scientific claims.
Question Answering
For this task, we consider the following datasets: 2WikiMultiHopQA [27], StrategyQA [28], MuSiQue [29], and HotpotQA [30]. To answer the questions in these datasets requires not only multi-hop reasoning but also cross-document analysis.
Reading Comprehension
Machine reading comprehension task requires a model to process documents and select an answer from the provided candidates to a question about the content [14]. We primarily consider WIKIHOP [31] in the task, which necessitates multi-hop reasoning to derive the final answer. Additionally, HotpotQA dataset [30] is also frequently considered as part of the question answering domain.
Appendix E Limitations
We propose an information re-organization approach to improve the reasoning of large language models, which performs well on some context-aware reasoning tasks but still has some limitations. Firstly, the structures of information re-organization are limited. Generally, there are multiple structures for information reorganization, such as tables, timelines, etc. Next, we will extend the re-organization structure to include more types. Secondly, the re-organization process relies on large language models. If we can implement this re-organization using smaller language models, our method will become more generalizable. This is another direction we need to focus on in the future.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Abstract, Introduction
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations in Appendix
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Experimental
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Supplementary material
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Experimental
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Supplementary material
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Conclusion
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Introduction, Related Work, Methodology, Experimental, Results and Analysis.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not involve the release of new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.